The machine element is subject to a moment = M = 75 N.m and a shear force V = 90 N. Determine the bending and shear stress at point C.20 mm 10 mm 101 20 mm 10 mm 10 mm B А 10 mm N V С M = 75 N·m 40 mm V=90 N
Answer #1

19 8ohm Tromm N A 10 40mm У 10 A = A2 = 40x10 = 400mm? 9,=Y₂ = 40 = 20mm A3 = 80x 10 = 800 mm Y3 = 40 + 10 = 45mm y : A9 14A2Nady bendoning stress at c m = 75Nm = 75x10mm 75mm 11.5mm Y = distance b/w cand NA 325 mm y = 7.5mm mxy 5x7.5x102 363-333 x 1